# Benchmarking Filtered Approximate Nearest Neighbor Search Algorithms on Transformer-based Embedding Vectors

### v1

@@ 4. A New Dataset to Benchmark FANNS
To address this gap, we introduce the arxiv-for-fanns dataset, where each item corresponds to a research paper. The dataset is available in three scales: small (1k items), medium (100k items), and large (over 2.7M items). Each item includes a 4096-dimensional, normalized text embedding of the paper abstract, computed using the stella_en_400M_v5 model (NovaSearch, 2025; Zhang et al., 2024), based on abstracts from the arXiv dataset (Cornell University et al., 2025), thus satisfying requirement 1. Among the top 10 models on the English MTEB leaderboard (Muennighoff et al., 2022), Stella is the most lightweight [^2] As of March 11, 2025 (Benchmark, 2025). Each item also contains 11 real-world attributes (see Table 4) covering all three attribute types (see Section 2.3), with diverse value distributions shown in Figure 14, fulfilling requirement 2. To satisfy requirement 3, each dataset variant includes three sets of 10k queries for EM, R, and EMIS filters (see Section 2.4), along with their corresponding precomputed ground truth. Query vectors are generated by embedding search terms synthesized by GPT-4o (Hurst et al., 2024), while filter values are sampled from the dataset attributes. Figure 15 illustrates the selectivity distribution (as defined in Section 3.1) across queries in arxiv-for-fanns-large. To support unfiltered ANNS evaluation, we also provide ground truth results without any filtering. All three versions of the dataset are publicly available on Hugging Face [^3] https://huggingface.co/datasets/SPCL/arxiv-for-fanns-large , thereby fulfilling requirement 4.
@@ 5. Benchmarking FANNS
We benchmark a selection of FANNS methods on our newly introduced arxiv-for-fanns dataset (see Section 4). Experiments with the medium-scale dataset are conducted on a laptop, while those on the large-scale dataset are executed on a compute cluster.
Method Experiment Parameters
medium, EM filter $M=16,M_{\beta}=24,\gamma=10$
medium, R filter $M=32,M_{\beta}=24,\gamma=12$
medium, EMIS filter $M=16,M_{\beta}=24,\gamma=15$
large, EM filter $M=32,M_{\beta}=32,\gamma=10$
large, R filter $M=32,M_{\beta}=16,\gamma=12$
large, EMIS filter $M=48,M_{\beta}=48,\gamma=15$
medium, EM filter $B=128$
large, EM filter $B=512$
medium, EM filter $R_{\text{small}}=32,L_{\text{small}}=80,R_{\text{stitched}}=48,\alpha=1.1$
medium, EMIS filter $R_{\text{small}}=32,L_{\text{small}}=100,R_{\text{stitched}}=48,\alpha=1.2$
large, EM filter $R_{\text{small}}=96,L_{\text{small}}=100,R_{\text{stitched}}=64,\alpha=1.4$
large, EMIS filter $R_{\text{small}}=64,L_{\text{small}}=80,R_{\text{stitched}}=64,\alpha=1.2$
medium, EM filter $K=100,L=80,S=15,R=200,\text{RANGE}=70,PL=200,B=0.5,\text{weight\_search}=1M$
large, EM filter $K=80,L=60,S=10,R=200,\text{RANGE}=60,PL=300,B=0.6,\text{weight\_search}=1M$
NHQ (nsw) (Wang et al., 2024) all $\text{efSearch}\in\{50,100,150,200,300,500,1k,2k,4k\}$
medium, EM filter $M=40,\text{MaxM0}=40,\text{efConstruction}=300,\text{weight\_search}=1M$
large, EM filter $M=40,\text{MaxM0}=60,\text{efConstruction}=150,\text{weight\_search}=1M$
medium, EM filter $\delta=2,R=24,L=80,\alpha=1.4,\sigma=16$
medium, EMIS filter $\delta=8,R=32,L=100,\alpha=1.4,\sigma=16$
large, EM filter $\delta=6,R=48,L=150,\alpha=1.6,\sigma=24$
large, EMIS filter $\delta=6,R=48,L=150,\alpha=1.4,\sigma=24$
[…]
@@ 5.1. Collected Metrics
We focus on the well-established recall vs. queries per second (QPS) plots, which illustrate the trade-off between accuracy and query throughput achieved by each method. We define recall@k as:
[…]
@@ 5.2. Parameter Search
Since all FANNS methods considered involve tunable parameters that influence index construction, we perform a dedicated parameter search for each combination of method, dataset scale (medium vs. large), and filter type prior to benchmarking. Given that some methods expose up to eight parameters, a full grid search is computationally infeasible. Instead, we adopt a greedy parameter search strategy, outlined in Algorithm 1. The algorithm takes as input a list of tunable parameters, their corresponding candidate values, and the index of each parameter’s default value inside the list of candidate values. Candidate and default values are selected based on the literature and publicly available implementations of each method. The get_reward() function performs two iterations of building the index with a given parameter configuration and querying it using 505050 randomly sampled queries from the dataset. It computes the recall vs. QPS curve averaged over both iterations and returns the highest QPS achieved at a recall of at least 0.95. If a method does not reach a recall of 0.95, it returns the highest achieved recall instead. The resulting parameter are listed in Table 5.
5.3. Benchmarking Methodology
Once the parameter search is complete, we benchmark each method five times (without warm-up) and report the mean and standard deviation for all measured metrics following scientific benchmarking practice (Hoefler and Belli, 2015). Index construction is performed using all available hardware threads on the respective system (details provided in the next paragraph), while query execution is restricted to a single thread [^4] Our script reports multiple threads during query execution of CAPS and SeRF, however, CPU utilization confirms that only one thread is active and the others are idle., following common practice (Cai et al., 2024; Gupta et al., 2023; Wang et al., 2024). We search for $k=10$ nearest neighbors and report recall@10, which is a standard choice in the literature (Wang et al., 2024; Yang et al., 2020c; Gollapudi et al., 2023; Patel et al., 2024; Cai et al., 2024). All benchmarks use Euclidean distance, as it is supported by all methods under evaluation. Since our transformer-based embeddings are normalized, KNN under Euclidean distance is equivalent to KNN under cosine distance. We exclude any preprocessing, such as sorting items by attribute value or transforming vectors into alternate formats when reporting index construction time.
@@ 5.3. Benchmarking Methodology
23:return {p : value_lists[p][best_indices[p]] for p in params}
@@ 5.4. Software and Hardware Configuration
To ensure reproducibility, we encapsulate our benchmarking infrastructure in a Docker container based on Ubuntu 20.04.6, which includes all necessary dependencies for the evaluated methods. Inside the container, we use Python 3.8.10, gcc 9.4.0, and g++ 9.4.0. Benchmarks on the arxiv-for-fanns-medium dataset are conducted on a laptop equipped with 16 GB of RAM and an Intel Core i7-1165G7 CPU with 4 physical cores and 8 hardware threads, running Arch Linux with kernel version 6.15.5-arch1-1. For the arxiv-for-fanns-large dataset, we execute benchmarks on a single node of a compute cluster with 384 GB of RAM and two Intel Xeon Gold 6154 CPUs, each with 18 physical cores and 36 hardware threads, running CentOS Linux 8.
@@ 5.5. Remarks on Algorithms
SeRF
Unlike all other methods, SeRF performs index construction and query execution within a single process, without persisting the index to disk. Therefore, we report only the overall peak memory usage, and we omit the index size from our results.
@@ 5.6. Results on arxiv-for-fanns-medium
Figure 17. 5.6) Index construction time (left), peak memory usage (middle), and index size (right) for the three filter types (EM, R, EMIS) on the arxiv-for-fanns-medium dataset.
Recall vs. QPS (Figure 16)
We observe that ACORN, the only evaluated method applicable to all three filter types, is generally outperformed by more specialized methods. For the EM filter, NHQ (kgraph), FDANN, and UNG achieve the best performance and perform similarly, whereas CAPS, NHQ (nsw), and ACORN are slower at comparable recall levels. The relative performance of ACORN, NHQ (kgraph), and UNG is consistent with previous benchmarks (Cai et al., 2024) on lower-dimensional datasets. FDANN, which showed inferior performance in those earlier studies, excels with our transformer-based embeddings, while NHQ (nsw) and CAPS appear to underperform. For the EMIS filter, previous benchmarks on lower-dimensional datasets (Cai et al., 2024) indicate that UNG outperforms FDANN and ACORN; however, on our transformer-based embeddings, FDANN demonstrates superior performance.
Index Construction Time (Figure 17, left)
We observe that index construction time varies by up to an order of magnitude across methods. CAPS is the fastest to build, followed by FDANN and NHQ (kgraph). SeRF, UNG, and NHQ (nsw) are significantly slower, with ACORN being the slowest across all filters.
Peak Memory Usage (Figure 17, middle)
Except for NHQ (kgraph) and SeRF, which show noticeably higher peak memory usage, differences in memory consumption during index construction and query execution are moderate across methods.
Index Size (Figure 17, right)
The index built by most methods is approximately 1.6 GB, corresponding to the size of the dataset. CAPS and NHQ (kgraph) produce very small indexes, only a few megabytes; however, they need to access the original database vectors during query execution. If we account for the size of the database vectors, the effective index size for these methods is also around 1.6 GB.
5.7. Challenges In Scaling Up the Dataset Size
[…]
Due to this large scale, 444 out of the 777 methods we benchmarked required modifications to their source code to run with the large dataset. UNG (Cai et al., 2024), CAPS (Gupta et al., 2023), and NHQ (KGraph) (Wang et al., 2024) all suffered from int32 overflows during memory allocation. Additionally, CAPS originally allocated a data structure on the stack, which we moved to the heap to support the large dataset. FDANN (Gollapudi et al., 2023), as mentioned in Section 5.5, initially stored two versions of the index, exceeding the local memory of our cluster’s compute node. Addressing the issues enabled benchmarking all 7 algorithms.

### v2

@@ 4. A New Dataset to Benchmark FANNS
Figure 19. 4) Index construction time, peak memory usage, and index size for EM filtering on four widely used datasets.
Query vectors are generated by embedding search terms synthesized by GPT-4o (Hurst et al., 2024), while filter values are sampled from the dataset attributes. Figure 21 illustrates the selectivity distribution (as defined in Section 3.1) across queries in arxiv-for-fanns-large. To support unfiltered ANNS evaluation, we also provide ground truth results without any filtering. All three versions of the dataset are publicly available on Hugging Face [^3] https://hf.co/datasets/SPCL/arxiv-for-fanns-small  [^4] https://hf.co/datasets/SPCL/arxiv-for-fanns-medium  [^5] https://hf.co/datasets/SPCL/arxiv-for-fanns-large , fulfilling criterion 3.
@@ 5. Benchmarking FANNS
We benchmark ACORN (Patel et al., 2024), CAPS (Gupta et al., 2023), two versions of FDANN (Gollapudi et al., 2023) (filtered and stitched), two versions of NHQ (Wang et al., 2024) (nsw and kgraph), SeRF (Zuo et al., 2024), iRangeGraph (Xu et al., 2024), DIGRA (Jiang et al., 2025), DSG (Peng et al., 2025), and UNG (Cai et al., 2024) on our arxiv-for-fanns dataset. This selection of methods covers a wide range of filtering approaches, indexing techniques, and filter types.
@@ 5.1. Collected Metrics
To illustrate the trade-off between accuracy and query throughput, we focus on recall vs. queries per second (QPS) plots as our primary evaluation metric. We define recall@k as:
$\text{recall@}k=\frac{|\text{knn}_{\text{alg}}\cap\text{knn}_{\text{gt}}|}{k}$
@@ 5.2. Parameter Search
Since all FANNS methods considered involve tunable parameters that influence index construction, we perform a dedicated parameter search for each combination of method, dataset, and filter type prior to benchmarking. Given that some methods expose up to eight parameters, a full grid search is computationally infeasible and we therefore rely on a greedy search strategy. For some combinations of method and dataset, even the greedy search takes multiple weeks to complete. Algorithm 1 presents the pseudocode of our parameter search algorithm.
@@
23:return {p : value_lists[p][indices[p]] for p in params}
This algorithm takes as input a list of tunable parameters, their corresponding candidate values, and the index of each parameter’s default value inside the list of candidate values. Candidate and default values are selected based on the literature and publicly available implementations of each method. The get_reward() function performs two iterations of building the index with a given parameter configuration and querying it using $50$ randomly sampled queries from the dataset. It computes the recall vs. QPS curve averaged over both iterations and returns the highest QPS achieved at a recall of at least 0.95. If a method does not reach a recall of 0.95, it returns the highest achieved recall instead.
To validate our parameter search strategy, we perform a full grid search for the ACORN algorithm (only 3 parameters) on the Audio dataset, which is much smaller than arxiv-for-fanns. Both search strategies found parameters $M=24$ and $\beta=32$. The greedy search settled for $\gamma=15$ while the grid search found $\gamma=12$. Figure 22 confirms that both parameter configurations yield nearly identical recall vs. QPS curves, and the different parameters are likely caused by randomness in index construction and query execution stages.
Figure 22. 3.1) Performance of ACORN on the Audio dataset with parameters from grid and greedy search.
The parameters identified by our search strategy for the SIFT, Audio, UQ-V, and Paper datasets used in Section 4 and for the arxiv-for-fanns-medium and arxiv-for-fanns-large datasets (see Sections 5.6 and 5.7) are reported in Table 4. For the study varying the parameter $k$ (see Section 5.8), we use the parameters obtained for $k=10$, as the choice of $k$ does not affect index construction. For the study on varying query selectivity (see Section 5.9), we perform a separate parameter search for each selectivity value. The resulting parameters are available in our repository [^6] https://github.com/spcl/fanns-benchmark .
5.3. Benchmarking Methodology
Once the parameter search is complete, we benchmark each method five times and report the mean and standard deviation for all measured metrics following scientific benchmarking practice (Hoefler and Belli, 2015). Index construction is performed using all available hardware threads on the respective system (see Section 5.4), while query execution is restricted to a single thread, following common practice (Cai et al., 2024; Gupta et al., 2023; Wang et al., 2024). Unless stated otherwise, we search for $k=10$ nearest neighbors and report recall@10, which is a standard choice in the literature (Wang et al., 2024; Yang et al., 2020c; Gollapudi et al., 2023; Patel et al., 2024; Cai et al., 2024). All benchmarks use Euclidean distance, as it is supported by all methods under evaluation. Since our transformer-based embeddings are normalized, KNN under Euclidean distance is equivalent to KNN under cosine distance. We exclude any preprocessing, such as sorting items by attribute value or transforming vectors into alternate formats when reporting index construction time to avoid making any assumptions on the format of the input data.
Method Experiment Parameters
SIFT, EM filter $M=32,M_{\beta}=16,\gamma=12$
Audio, EM filter $M=24,M_{\beta}=32,\gamma=12$
UQ-V, EM filter $M=24,M_{\beta}=32,\gamma=12$
Paper, EM filter $M=32,M_{\beta}=24,\gamma=12$
arxiv-medium, EM filter $M=16,M_{\beta}=24,\gamma=10$
arxiv-medium, R filter $M=32,M_{\beta}=24,\gamma=12$
arxiv-medium, EMIS filter $M=16,M_{\beta}=24,\gamma=15$
arxiv-large, EM filter $M=32,M_{\beta}=32,\gamma=10$
arxiv-large, R filter $M=32,M_{\beta}=16,\gamma=12$
arxiv-large, EMIS filter $M=48,M_{\beta}=48,\gamma=15$
SIFT, EM filter $B=256$
Audio, EM filter $B=256$
UQ-V, EM filter $B=128$
Paper, EM filter $B=1024$
arxiv-medium, EM filter $B=128$
arxiv-large, EM filter $B=512$
SIFT, EM filter $R_{\text{small}}=32,L_{\text{small}}=80,R_{\text{stitched}}=48,\alpha=1.3$
Audio, EM filter $R_{\text{small}}=16,L_{\text{small}}=80,R_{\text{stitched}}=96,\alpha=1.2$
UQ-V, EM filter $R_{\text{small}}=32,L_{\text{small}}=80,R_{\text{stitched}}=64,\alpha=1.1$
Paper, EM filter $R_{\text{small}}=32,L_{\text{small}}=100,R_{\text{stitched}}=128,\alpha=1.3$
arxiv-medium, EM filter $R_{\text{small}}=32,L_{\text{small}}=80,R_{\text{stitched}}=48,\alpha=1.1$
arxiv-medium, EMIS filter $R_{\text{small}}=32,L_{\text{small}}=100,R_{\text{stitched}}=48,\alpha=1.2$
arxiv-large, EM filter $R_{\text{small}}=96,L_{\text{small}}=100,R_{\text{stitched}}=64,\alpha=1.4$
arxiv-large, EMIS filter $R_{\text{small}}=64,L_{\text{small}}=80,R_{\text{stitched}}=64,\alpha=1.2$
FDANN (filtered) (Gollapudi et al., 2023) all $L_{\text{search}}\in\{10,20,30,50,100,150,200,300,500,1k\}$
SIFT, EM filter $R=48,L=100,\alpha=1.0$
Audio, EM filter $R=32,L=100,\alpha=1.2$
UQ-V, EM filter $R=96,L=60,\alpha=1.0$
Paper, EM filter $R=48,L=100,\alpha=1.3$
arxiv-medium, EM filter $R=48,L=100,\alpha=1.0$
arxiv-medium, EMIS filter $R=48,L=80,\alpha=1.2$
arxiv-large, EM filter $R=48,L=200,\alpha=1.2$
arxiv-large, EMIS filter $R=48,L=150,\alpha=1.2$
SIFT, EM filter $K=100,L=100,S=15,R=300,\text{RANGE}=50,PL=400,B=0.4,\text{weight}=1M$
Audio, EM filter $K=40,L=40,S=4,R=300,\text{RANGE}=70,PL=400,B=0.6,\text{weight}=10B$
UQ-V, EM filter $K=80,L=80,S=10,R=200,\text{RANGE}=60,PL=300,B=0.5,\text{weight}=1M$
Paper, EM filter $K=60,L=40,S=15,R=400,\text{RANGE}=20,PL=200,B=0.3,\text{weight}=100M$
arxiv-medium, EM filter $K=100,L=80,S=15,R=200,\text{RANGE}=70,PL=200,B=0.5,\text{weight}=1M$
arxiv-large, EM filter $K=80,L=60,S=10,R=200,\text{RANGE}=60,PL=300,B=0.6,\text{weight}=1M$
NHQ (nsw) (Wang et al., 2024) all $\text{ef\_search}\in\{50,100,150,200,300,500,1k,2k,4k\}$
SIFT, EM filter $M=50,\text{MaxM0}=50,\text{efConstruction}=200,\text{weight\_search}=1M$
Audio, EM filter $M=40,\text{MaxM0}=50,\text{efConstruction}=300,\text{weight\_search}=1M$
UQ-V, EM filter $M=20,\text{MaxM0}=40,\text{efConstruction}=300,\text{weight\_search}=1M$
Paper, EM filter $M=40,\text{MaxM0}=50,\text{efConstruction}=300,\text{weight\_search}=1M$
arxiv-medium, EM filter $M=40,\text{MaxM0}=40,\text{efConstruction}=300,\text{weight\_search}=1M$
arxiv-large, EM filter $M=40,\text{MaxM0}=60,\text{efConstruction}=150,\text{weight\_search}=1M$
SIFT, EM filter $\delta=6,R=32,L=100,\alpha=1.4,\sigma=12$
Audio, EM filter $\delta=6,R=48,L=100,\alpha=1.4,\sigma=16$
UQ-V, EM filter $\delta=6,R=48,L=100,\alpha=1.2,\sigma=16$
Paper, EM filter $\delta=6,R=32,L=100,\alpha=1.4,\sigma=16$
arxiv-medium, EM filter $\delta=2,R=24,L=80,\alpha=1.4,\sigma=16$
arxiv-medium, EMIS filter $\delta=8,R=32,L=100,\alpha=1.4,\sigma=16$
arxiv-large, EM filter $\delta=6,R=48,L=150,\alpha=1.6,\sigma=24$
arxiv-large, EMIS filter $\delta=6,R=48,L=150,\alpha=1.4,\sigma=24$
SeRF (Zuo et al., 2024) all $\text{ef\_search}\in\{4,8,16,32,64,128,256,512,1024\}$
arxiv-medium, R filter $\text{index\_k}=100,\text{ef\_construction}=100,\text{ef\_max}=500$
arxiv-large, R filter $\text{index\_k}=100,\text{ef\_construction}=100,\text{ef\_max}=600$
iRangeGraph (Xu et al., 2024) all $\text{ef\_search}\in\{1,2,3,4,5,6,8,10,15,20,30,50,100,200\}$
arxiv-medium, R filter $M=32,\text{ef\_construction}=200$
arxiv-large, R filter $M=24,\text{ef\_construction}=100$
DIGRA (Jiang et al., 2025) arxiv-medium, R filter $\text{ef\_search}\in\{1,2,3,4,5,6,8,10,15,20,30,50,100,200\},M=24,\text{ef\_construction}=200$
DSG (Peng et al., 2025) arxiv-medium, R filter $\text{ef\_search}\in\{4,8,16,32,64,128,256,512,1024\},M=24,\text{ef\_construction}=200,\text{ef\_max}=500,\alpha=1.0$
@@ 5.4. Software and Hardware Configuration
We encapsulate our benchmarking infrastructure in a Docker container based on Ubuntu 20.04.6. We use Python 3.8.10, gcc 9.4.0, and g++ 9.4.0. Benchmarks on the medium-scale dataset are run on an Intel Core i7-1165G7 CPU with 4 physical cores, 8 hardware threads, and 16 GB RAM, running Arch Linux with kernel version 6.15.5-arch1-1 and those on the large dataset run on a machine with 384 GB of RAM and two Intel Xeon Gold 6154 CPUs, each with 18 physical cores and 36 hardware threads, running CentOS Linux 8.
Figure 24. 5.6) Index construction time, peak memory usage, and index size on the arxiv-for-fanns-medium dataset.
@@ 5.5. Remarks on Algorithms
SeRF, DIGRA
Unlike most other methods, SeRF and DIGRA perform index construction and query execution within a single process, without persisting the index to disk. Therefore, we report only the overall peak memory usage, and we omit the index size from our results.
@@
DSG, DIGRA
Unlike most of the other methods, DSG and DIGRA do not parallelize index construction. Therefore, we report index construction time for a single thread in our benchmarking on arxiv-for-fanns-medium and we omit DSG and DIGRA from our benchmarking on arxiv-for-fanns-large.
@@ 5.6. Results on arxiv-for-fanns-medium
Figure 23 shows the recall vs. QPS curves for the three filter types on the arxiv-for-fanns-medium dataset. We observe that ACORN, the only evaluated method applicable to all three filter types, is generally outperformed by more specialized methods. FDANN, which lagged slightly behind UNG on the established datasets (see Figure 18), matches UNG’s performance on our transformer-based embeddings with EM filters and even outperforms UNG with EMIS filters. NHQ (kgraph), a top performer on only two of the four established datasets, is among the best methods on our dataset.
Observation [Uncaptioned image]
Specialized FANNS methods that only support specific filters outperform more versatile methods by up to $10\times$.
Figure 24 summarizes index construction time, peak memory usage, and index size on the arxiv-for-fanns-medium dataset. Index construction time varies by more than an order of magnitude across methods, with DIGRA and DSG being the slowest since they are the only methods that do not parallelize index construction. Except for NHQ (kgraph) and SeRF, which show noticeably higher peak memory usage, differences in memory consumption during index construction and query execution are moderate across methods. The index built by most methods is approximately 1.6 GB, corresponding to the size of the dataset. Some methods produce very small indexes but require access to the original database vectors during query execution, resulting in the same effective index size.
Observation [Uncaptioned image]
Index construction times vary by up to $16\times$ across methods but do not correlate with query performance.
Figure 25. 5.7) Recall@10 vs. QPS plots for the three filter types (EM, R, EMIS) on the arxiv-for-fanns-large dataset.
Figure 26. 5.7) Index construction time, peak memory usage, and index size on the arxiv-for-fanns-large dataset.
@@ 5.7. Results on arxiv-for-fanns-large
5.7.1. Challenges In Scaling Up the Dataset Size
While $2.7$ million vectors may not sound particularly large, note that our transformer-based embeddings have 4,0964{,}096 dimensions, making our dataset over $85\times$ larger than the well-known SIFT1M dataset (Laurent Amsaleg, 2025) which uses 128-dimensional embedding vectors. We had to modify the source code of four out of the eleven methods we benchmarked to fix int32 overflows during memory allocation (UNG (Cai et al., 2024), CAPS (Gupta et al., 2023), and NHQ (KGraph) (Wang et al., 2024)) or to reduce the index size (FDANN (stitched)  (Gollapudi et al., 2023)) to fit in the local persistent storage. Because index construction for DSG and DIGRA is non-parallelized, takes over a day per run, and requires dozens of runs for parameter tuning, we could not benchmark these two methods on the large dataset.
5.7.2. Results
Figure 25 shows the recall vs. QPS curves for the three filter types on the arxiv-for-fanns-large dataset. Comparing these results to those on the medium-scale dataset in Figure 23 provides insights into scalability. ACORN, both NHQ variants, FDANN (filtered), SeRF, and iRangeGraph maintain their relative performance, with throughput dropping by only a factor of about $4$ despite the dataset being $27\times$ larger. This indicates good scalability, especially considering that both experiments used a single thread, albeit on different machines. FDANN (stitched), CAPS, and UNG perform poorly on the large dataset and fail to exceed $25\%$ recall. We cannot rule out that suboptimal parameter choices contribute to this performance drop as our greedy parameter search may have converged to local optima. This highlights the difficulty of parameter tuning for FANNS methods on large datasets.
Observation [Uncaptioned image]
Scaling to large datasets can cause recall to drop below $0.2$, rendering some methods useless. Adjustments of the implementation or parameter tuning strategies may be required.
Figure 26 reports index construction time, peak memory usage, and index size for all methods. On the $27\times$ larger dataset, construction time increased by roughly an order of magnitude. Since $9\times$ more threads were used, this suggests super-linear growth with dataset size or sublinear thread speedup. CAPS remains the fastest to build, with NHQ and FDANN also being relatively fast. UNG has by far the longest construction time, likely because it failed to reach the $0.95$ recall target and the search shifted toward more expensive configurations. Memory usage scales roughly linearly with dataset size, and relative differences remain similar to the medium-scale case. UNG is notable for compressing the index to about $25\%$ of the dataset with CAPS, iRangeGraph, and NHQ (kgraph) producing compact indexes but requiring access to the original vectors.
Observation [Uncaptioned image]
Parameter tuning is difficult on large datasets, with indexes construction taking up to 6 hours for 2.7M items.
Figure 27. 5.8) Recall@k vs. QPS plots for varying values of $k$ on the arxiv-for-fanns-medium dataset with EM filtering.
Figure 28. 5.8) Recall@k vs. QPS plots for varying values of $k$ on the arxiv-for-fanns-medium dataset with R filtering.